# Faithful Group Shapley Value

### v1

@@ 1 Introduction
As data become increasingly crucial in modern machine learning, quantifying its value has significant implications for fair compensation and data market design [29, 39]. The Shapley value, a foundational concept from cooperative game theory [30], stands out as the unique valuation method satisfying four desirable axioms for fair data valuation [6, 13]. Consequently, it has been applied to diverse applications, including collaborative intelligence [22, 25] and copyright compensation [35], and quantifying feature importance in explainable AI [26].
@@
Many real-world scenarios demand evaluation of data sets instead of individual data points. In applications like data marketplaces [39] or compensation allocation for generative AI [4, 35], data are contributed by owners who possess entire datasets, making group-level data valuation a natural choice. A more technical motivation is that individual-level Shapley values are often computationally challenging to approximate for big data; while group-level valuation is much faster and can provide useful insights [16]. In explainable AI, group-level feature importance reportedly provides more robust and interpretable valuation [32, 3].
Prior works extend the Shapley framework to group data valuation by treating each pre-defined group as an atomic unit and then applying the standard Shapley value formulation [14, 20, 35, 32], which we refer to as Group Shapley Value (GSV). However, we demonstrate—both theoretically and empirically—that this approach suffers from susceptibility to strategic manipulation through partitioning. Consider the valuation of a fixed data group $A$. As shown in Figure 1, partitioning the remaining data points into smaller subgroups reduces the GSV for $A$, despite no change in data content. In other words, unfair players may exploit the loophole in GSV by splitting their data among puppy subsidiaries, a vulnerability we coin as shell company attack.
@@
While FGSV safeguards fairness, its exact computation requires combinatorial computation of individual Shapley values. Although numerous approximation algorithms for individual Shapley values have been developed [2, 6, 12, 34, 26, 3, 38, 24, 27], simply applying them and summing up results in each group is slow for large groups and may compound approximation errors. Importantly, we proved that a small subset of terms dominate in FGSV’s formula. As a result, we developed an original algorithm that directly approximates FGSV fast and accurately. Our numerical experiments demonstrate its computational efficiency and approximation accuracy against summing up individual Shapley values computed by state-of-the-art (SOTA) methods. We also applied FGSV to fair copyright compensation on Stable Diffusion models for image generation.
@@ 2 Preliminaries: individual and group Shapley values
Individual Shapley value (SV). Let $\mathcal{D}=\{z_{1},\ldots,z_{n}\}\in\mathcal{Z}^{n}$ denote the training data, where $n:=|\mathcal{D}|$. Let $S\subseteq[n]$ be an index set of cardinality $s:=|S|$, and write ${\cal S}:=\{z_{i}:i\in S\}$. A utility function $U({\cal S})$ assigns a performance score (such as classification accuracy) to method trained based on data ${\cal S}$. To simplify notation, we may refer to “$U({\cal S})$” by index as “$U(S)$” and write them interchangeably. The individual Shapley value of data point $i$ is defined as
@@
Group Shapley value (GSV). As mentioned in Section 2, most group-level Shapley values [14, 35, 32] adopt the so-called group-as-individual (GaI) approach. Suppose the entire data set is partitioned into $(K+1)$ disjoint groups: $\mathcal{D}=S_{0}\cup S_{1}\cup\cdots\cup S_{K}$, in which, $S_{0}\subseteq[n]$ is the group we aim to evaluate. The group Shapley value (GSV) of $S_{0}$, as in [14, 35, 32], is defined as
@@
Suppose in $\mathcal{D}=\{z_{1},\dots,z_{n}\}$ are an i.i.d. sample. Let $\bar{U}(s)=\mathbb{E}_{S\sim\mathcal{P}^{s}}[U(S)]$ denote the expected utility for data $S$ (which only depends on $s=|S|$). Now we split a group $S_{k}$ into two non-empty subgroups $S_{k}^{\prime}$ and $S_{k}^{\prime\prime}$ (i.e., $S_{k}^{\prime}\cup S_{k}^{\prime\prime}=S_{k}$, $S_{k}^{\prime}\cap S_{k}^{\prime\prime}=\emptyset$, and $S_{k}^{\prime},S_{k}^{\prime\prime}\neq\emptyset$), If $\bar{U}(s)$ satisfies a prudence condition:
@@
Condition (3) is a familiar concept in economics, characterizing risk-aversion behaviors under uncertainty [7]. It is also observed in machine learning, where the performance of a learning method saturates as the same type of data repeatedly come in [17, 10].
@@ 3.2 Fast and accurate approximation algorithm for FGSV
Assumption 1 is a common regularity condition, and Assumption 2 introduces a mild second-order stability requirement. In Section 3.5, we will show examples that Assumption 2 is satisfied by some commonly used utility functions.
@@
However, estimating $\mu(\frac{s_{1}+1}{s};s,s_{0},n)$ and $\mu(\frac{s_{1}}{s};s,s_{0},n)$ separately can be statistically inefficient, as the noise from two independent Monte Carlo estimates could mask the signal in their difference—especially when the true gap is small for large $s$. Following the variance-reduction technique used in stochastic simulation [21], we propose to estimate $\Delta\mu$ directly using paired Monte Carlo terms:
@@
where the tuple $\{(S^{(j)},i_{1}^{(j)},i_{2}^{(j)})\}$ is i.i.d. sampled from $\big{\{}(S,i_{1},i_{2}):|S|=s-1,|S\cap S_{0}|=s_{1},i_{1}\in S_{0}\setminus S,i_{2}\in S_{0}^{c}\setminus S\big{\}}$.
@@ 3.3 Computational complexity
Our theoretical analysis for approximating $\mathcal{T}(s)$ relies on the stability of the utility function $U$. Specifically, we employ the concept of deletion stability to quantify the maximum change in the utility function when a single data point is removed [1, 9].
@@
Suppose Assumptions 1 and 2 hold, and $U$ is $\beta(s)$-deletion stable with $\beta(s)=O(s^{-1})$. Let $\alpha_{0}=s_{0}/n$. For any $\epsilon,\delta\in(0,1)$, set $\bar{s}\asymp\epsilon^{-1/\upsilon}$, $m_{1}\asymp\epsilon^{-(4+2\upsilon)/\upsilon}\log(n/\delta)$, and $m_{2}\asymp\max\{1,\epsilon^{-2}(\alpha_{0}(1-\alpha_{0}))^{2}\log^{3}(n/\delta)\}$. Algorithm 1 gives an $(\epsilon,\delta)$-approximation of $\operatorname{FGSV}(S_{0})$ with $O\left(n\left[1+(\alpha_{0}(1-\alpha_{0}))^{2}(\log n)^{3}\right]\right)$ utility evaluations.
The assumption of $\beta(s)=O(1/s)$ deletion stability in Theorem 3 is mild. Comparable assumptions have appeared in the literature of individual Shapley [19, 36, 37] and other learning problems [1, 9].
Numerous recent advances have developed efficient algorithms for approximating individual Shapley values [13, 34, 24, 26, 3, 27, 23, 38]. These methods target an $(\epsilon,\delta)$-approximation on the full individual Shapley-value vector $\|\widehat{\theta}-\theta\|_{2}$, where $\theta:=(\operatorname{SV}(1),\ldots,\operatorname{SV}(n))$ and $\widehat{\theta}$ is the approximation for $\theta$. The SOTA method achieves this guarantee in $O(n\epsilon^{-2}\log(n/\delta))$ utility evaluations. Approximating $\operatorname{FGSV}(S_{0})=\sum_{i\in S_{0}}\theta_{i}$ by simply summing up individual approximations can lead to an additive error bounded by $\sqrt{s_{0}}\|\widehat{\theta}-\theta\|_{2}$. Thus, achieving the same $(\epsilon,\delta)$-approximation on $\operatorname{FGSV}(S_{0})$ requires tightening the error tolerance to $\epsilon/\sqrt{s_{0}}$, increasing the sample complexity to $O\big{(}\alpha_{0}n^{2}\mathrm{Poly}(\log n)\big{)}$. By directly targeting the group objective, our method avoids this quadratic blow-up and offers a significant speed-up over the SOTA.
@@ 3.4 Utility function values for small input sets
Our remedy is very simple. We set a threshold for input size, denoted by $B$, such that $U(S)$ is always well-defined for $|S|\geq B$. If $|S|<B$, we inject $B-|S|$ non-informative data points, to elevate the input size to $B$; otherwise no change is made to $U(S)$. Clearly, we can expect that as the size of informative data (i.e., $|S|$) increases, less amount of non-informative data will be injected, and $U(S)$ gradually becomes non-baseline. For more implementation details, see Algorithm 2 in Appendix B.
@@ 3.5.1 Stochastic Gradient Descent (SGD)
3.5.2 Influence Function (IF)
The influence function (IF) method [8, 18] considers the following regularized empirical loss:
Like in Section 3.5.1, suppose the utility can be written as $U(S)=u(\widehat{\theta}_{S})$. Under mild conditions (see Proposition 3), standard IF theory [8, 18] implies that
for any $z_{1},z_{1}^{\prime}$, where $H_{\widehat{\theta}_{\mathcal{S}}}$ is the Hessian of ${\cal L}$ evaluated at $\widehat{\theta}_{\mathcal{S}}$. This yields the following proposition.
@@ 3.5.2 Influence Function (IF)
Suppose $u$ is continuously differentiable with $L$-Lipschitz and bounded gradient. Assume that the loss function $\ell(\theta;z)$ is convex in $\theta$, three times continuously differentiable with all derivatives up to order three uniformly bounded. Then for any $z_{1},z_{1}^{\prime},z_{2},z_{2}^{\prime}\in\mathcal{Z}$,
@@ 4 Experiments
We empirically compare our method (FGSV) to the approach of first computing individual Shapley values using SOTA methods and then summing them to form a group valuation. The results demonstrate the superiority of our method in both speed and approximation accuracy and its faithfulness under shell company attack in applications such as copyright attribution in generative AI and explainable AI. We report the key findings in the main paper and relegate full details to Appendix A.
@@ 4.1 Approximation accuracy and computational efficiency
In this experiment, we compare our method to the following benchmarks: (1) Permutation-based estimator [2, 6] that averages marginal contributions over random data permutations; (2) Group Testing-based estimator [12, 34] that uses randomized group inclusion tests to estimate pairwise Shapley value differences; (3) Complementary Contribution estimator [38] that uses stratified sampling of complementary coalitions; (4) One-for-All estimator [24] that uses weighted subsets and reuses utility evaluations; (5) KernelSHAP [26], a regression-based method with locally weighted samples; (6) Unbiased KernelSHAP [3], KernelSHAP with a bias-correction; and (7) LeverageSHAP [27] that speeds up KernelSHAP by weighted sampling based on leverage scores.
@@
We consider the Sum-of-Unanimity (SOU) cooperative game [24], where the individual Shapley values have a closed-form expression. Specifically, the utility function is defined as $U(S)=\sum_{j=1}^{d}\alpha_{j}\mathbf{1}_{\mathcal{A}_{j}\subseteq S}$, and the corresponding Shapley value for player $i\in[n]$ is given by $\operatorname{SV}(i)=\sum_{j=1}^{d}\frac{\alpha_{j}}{|\mathcal{A}_{j}|}\mathbf{1}_{i\in\mathcal{A}_{j}}$. Here, each subset $\mathcal{A}_{j}$ is generated by sampling a random size uniformly from $1$ to $n$ and then drawing that many players without replacement. We set $d=n^{2}$ and the coefficient $\alpha_{j}$ as the average of the weights of all players $i\in\mathcal{A}_{j}$, where each player $i$ has a weight $\frac{(i\bmod 4)}{4}$. We partition all players into $4$ groups based on their indices: $S_{k}=\big{\{}i\in[n]:i\equiv k-1(\bmod 4)\big{\}}$ for $k\in[4]$, and our goal is to estimate the group value $\operatorname{FGSV}(S_{k})=\sum_{i\in S_{k}}\operatorname{SV}(i)$. Given a fixed budget of 20,000 utility function evaluations, we record the absolute relative error of the $\operatorname{FGSV}$ estimate every 200 iterations for each method. We summarize the performance using the Area Under the Convergence Curve (AUCC): $\operatorname{AUCC}(S_{k})=\frac{1}{100}\sum_{t=1}^{100}|\frac{\operatorname{FGSV}(S_{k})-\widehat{\operatorname{FGSV}}^{(200\cdot t)}(S_{k})}{\operatorname{FGSV}(S_{k})}|,$ where $\widehat{\operatorname{FGSV}}^{(\tau)}(\cdot)$ is the estimate after $\tau$ iterations. Lower AUCC indicates better accuracy and faster convergence. We also record the average runtime per iteration for each method.
Figure 2: Performance comparison in the SOU game. Top: Our method (FGSV) achieves the lowest AUCC across all problem sizes. Bottom: Our method costs the lowest runtime per iteration.

### v2

@@ 1 Introduction
As data become increasingly crucial in modern machine learning, quantifying its value has significant implications for faithful compensation and data market design (Sestino et al., 2025; Zhang et al., 2024). The Shapley value, a foundational concept from cooperative game theory (Shapley, 1953), stands out as the unique valuation method satisfying four desirable axioms for faithful data valuation (Ghorbani and Zou, 2019; Jia et al., 2019). Consequently, it has been applied to diverse applications, including collaborative intelligence (Li et al., 2021; Liu et al., 2022) and copyright compensation (Wang et al., 2024), and quantifying feature importance in explainable AI (Lundberg, 2017).
@@
Many real-world scenarios demand evaluation of data sets instead of individual data points. In applications like data marketplaces (Zhang et al., 2024) or compensation allocation for generative AI (Deng and Ma, 2023; Wang et al., 2024), data are contributed by owners who possess entire datasets, making group-level data valuation a natural choice. A more technical motivation is that individual-level Shapley values are often computationally challenging to approximate for big data; while group-level valuation is much faster and can provide useful insights (Kang et al., 2024). In explainable AI, group-level feature importance reportedly provides more robust and interpretable valuation (Wang et al., 2025; Covert and Lee, 2021).
As group data valuation frameworks are increasingly used in real-world scenarios such as data markets and copyright compensation, their robustness against adversarial manipulation becomes important. Prior studies on attacks and defenses in data valuation have focused primarily on the individual-level data valuation, such as the copier attack that duplicates existing data to unjustly inflate value (Falconer et al., 2025; Han et al., 2022). In contrast, we identify a new vulnerability unique to group data valuation, which has been largely unexplored in literature. Existing extensions of the Shapley framework to group data valuation treat each pre-defined group as an atomic unit and then apply the standard Shapley value formulation (Jullum et al., 2021; Kwon et al., 2024; Wang et al., 2024, 2025), which we refer to as Group Shapley Value (GSV). However, we demonstrate—both theoretically and empirically—that this approach suffers from susceptibility to strategic manipulation through partitioning. Consider the valuation of a fixed data group $A$. As shown in Figure 1, partitioning the remaining data points into smaller subgroups reduces the GSV for $A$, despite no change in data content. In other words, malicious players may exploit the loophole in GSV by splitting their data among puppy subsidiaries, a vulnerability we coin as shell company attack.
@@
While FGSV safeguards faithfulness, its exact computation requires combinatorial computation of individual Shapley values. Although numerous approximation algorithms for individual Shapley values have been developed (Castro et al., 2009; Ghorbani and Zou, 2019; Jia et al., 2019; Wang and Jia, 2023a; Lundberg, 2017; Covert and Lee, 2021; Zhang et al., 2023; Li and Yu, 2024b; Musco and Witter, 2025), simply applying them and summing up results in each group is slow for large groups and may compound approximation errors. Importantly, we proved that a small subset of terms dominate in FGSV’s formula. As a result, we developed an original algorithm that directly approximates FGSV fast and accurately. Our numerical experiments demonstrate its computational efficiency and approximation accuracy against summing up individual Shapley values computed by state-of-the-art (SOTA) methods. We also applied FGSV to faithful copyright compensation on Stable Diffusion models for image generation.
@@ 2 Preliminaries: individual and group Shapley values
Individual Shapley value (SV). Let $\mathcal{D}=\{z_{1},\ldots,z_{n}\}\in\mathcal{Z}^{n}$ denote the training data, where $n:=|\mathcal{D}|$. Let $S\subseteq[n]$ be an index set of cardinality $s:=|S|$, and write ${\cal S}:=\{z_{i}:i\in S\}$. A utility function $U({\cal S})$ assigns a performance score (such as classification accuracy) to method trained based on data $\cal S$}. To simplify notation, we may refer to $U({\cal S})$” by index as “U(S)U(S)” and write them interchangeably. The individual Shapley value of data point $i$ is defined as
@@
Group Shapley value (GSV). As mentioned in Section 2, most group-level Shapley values (Jullum et al., 2021; Wang et al., 2024, 2025) adopt the so-called group-as-individual (GaI) approach. Suppose the entire data set is partitioned into $(K+1)$ disjoint groups: $\mathcal{D}=S_{0}\cup S_{1}\cup\cdots\cup S_{K}$, in which, $S_{0}\subseteq[n]$ is the group we aim to evaluate. The group Shapley value (GSV) of $S_{0}$, as in Jullum et al. (2021); Wang et al. (2024, 2025), is defined as
@@
Let $\mathcal{P}$ be the underlying data distribution over the sample space $\mathcal{Z}$, and write $\mathcal{P}^{s}$ for the $s$-fold product distribution; thus $S\sim\mathcal{P}^{s}$ denotes an i.i.d. sample $S=(z_{1},\ldots,z_{s})\in\mathcal{Z}^{s}$ with $z_{i}\sim\mathcal{P}$. Let $\bar{U}(s)=\mathbb{E}_{S\sim\mathcal{P}^{s}}[U(S)]$ denote the expected utility for data $S$ (which only depends on $s=|S|$). Now we split a group $S_{k}$ into two non-empty subgroups $S_{k}^{\prime}$ and $S_{k}^{\prime\prime}$ (i.e., $S_{k}^{\prime}\cup S_{k}^{\prime\prime}=S_{k}$, $S_{k}^{\prime}\cap S_{k}^{\prime\prime}=\emptyset$, and $S_{k}^{\prime},S_{k}^{\prime\prime}\neq\emptyset$), If $\bar{U}(s)$ satisfies a prudence condition:
@@
Condition (3) is a familiar concept in economics, characterizing risk-aversion behaviors under uncertainty (Gollier, 2001). It is also observed in machine learning, where the performance of a learning method saturates as the same type of data repeatedly come in (Kaplan et al., 2020; Hestness et al., 2017).
@@ 3.2 Fast and accurate approximation algorithm for FGSV
Assumption 1 is a standard regularity condition commonly adopted in the data valuation literature (Jia et al., 2019; Wang and Jia, 2023b), and Assumption 2 introduces a mild second-order stability requirement. In Section 3.5, we will show examples that Assumption 2 is satisfied by some commonly used utility functions.
@@
However, estimating $\mu(\frac{s_{1}+1}{s};s,s_{0},n)$ and $\mu(\frac{s_{1}}{s};s,s_{0},n)$ separately can be statistically inefficient, as the noise from two independent Monte Carlo estimates could mask the signal in their difference—especially when the true gap is small for large $s$. Following the variance-reduction technique used in stochastic simulation (Law and Kelton, 1982), we propose to estimate $\Delta\mu$ directly using paired Monte Carlo terms:
@@
where the tuple $\{(S^{(j)},i_{1}^{(j)},i_{2}^{(j)})\}$ is i.i.d. sampled from $\big\{(S,i_{1},i_{2}):|S|=s,|S\cap S_{0}|=s_{1},i_{1}\in S_{0}\setminus S,i_{2}\in S_{0}^{c}\setminus S\big\}$.
@@ 3.3 Computational complexity
Our theoretical analysis for approximating $\mathcal{T}(s)$ relies on the stability of the utility function $U$. Specifically, we employ the concept of deletion stability to quantify the maximum change in the utility function when a single data point is removed (Bousquet and Elisseeff, 2002; Hardt et al., 2016).
@@
The regime $\beta(s)=O(1/s)$ for deletion stability is commonly assumed in the literature on individual Data Shapley approximation (Wu et al., 2023, 2024) and in the analysis of algorithm stability (Bousquet and Elisseeff, 2002; Hardt et al., 2016).
@@
Suppose the utility function $U$ is $O(1/s)$-deletion stable, then Algorithm 1 guarantees that for any truncation threshold $\bar{s}$ and sample sizes $m_{1},m_{2}$, with probability at least $1-\delta$,
$\bigl|\widehat{\mathrm{FGSV}}(S_{0})-\mathrm{FGSV}(S_{0})\bigr|\lesssim\bar{s}\sqrt{\frac{\log(n/\delta)}{m_{1}}}+\alpha_{0}(1-\alpha_{0})\sqrt{\frac{\log(n/\delta)}{m_{2}}}\log n+\alpha_{0}(1-\alpha_{0})\bar{s}^{-\upsilon}.$
Specifically, choosing
$\bar{s}\asymp\epsilon^{-1/\upsilon},\quad m_{1}\asymp\epsilon^{-\frac{4+2\upsilon}{\upsilon}}\log\bigl(n/\delta\bigr),\quad m_{2}\asymp\max\Bigl\{{1,\epsilon^{-2}(\alpha_{0}(1-\alpha_{0}))^{2}(\log(n/\delta))^{3}\Bigr\}},$
yields an $(\epsilon,\delta)$‑approximation of $\mathrm{FGSV}(S_{0})$ with $O\left(n\cdot\max\left\{1,(\alpha_{0}(1-\alpha_{0}))^{2}(\log n)^{3}\right\}\right)$ utility evaluations.
Theorem 3 implies that our algorithm requires only $O(n\,\mathrm{Poly}(\log n))$ utility evaluations to achieve an $(\epsilon,\delta)$-approximation for the FGSV of a group $S_{0}$ whose size scales as a constant fraction of $n$.
To compare our method’s computational complexity with existing works, we notice numerous recent efficient algorithms for approximating individual Shapley values (Jia et al., 2019; Wang and Jia, 2023a; Li and Yu, 2024b; Lundberg, 2017; Covert and Lee, 2021; Musco and Witter, 2025; Li and Yu, 2024a; Zhang et al., 2023). They target an $(\epsilon,\delta)$-approximation on the full individual Shapley-value vector $\|\widehat{\theta}-\theta\|_{2}$, where $\theta:=(\operatorname{SV}(1),\ldots,\operatorname{SV}(n))$ and $\widehat{\theta}$ is the approximation for $\theta$. The SOTA method achieves this guarantee in $O(n\epsilon^{-2}\log(n/\delta))$ utility evaluations. Approximating $\operatorname{FGSV}(S_{0})=\sum_{i\in S_{0}}\theta_{i}$ by simply summing up individual approximations can lead to an additive error bounded by $\sqrt{s_{0}}\|\widehat{\theta}-\theta\|_{2}$. Thus, achieving the same $(\epsilon,\delta)$-approximation on $\operatorname{FGSV}(S_{0})$ requires tightening the error tolerance to $\epsilon/\sqrt{s_{0}}$, increasing the sample complexity to $O\big(\alpha_{0}n^{2}\mathrm{Poly}(\log n)\big)$. By directly targeting the group objective, our method avoids this quadratic blow-up and offers a significant speed-up over the SOTA. It continues to outperform SOTA individual-based methods even when evaluating multiple groups in parallel, provided that the number of groups is $o(n)$.
@@ 3.4 Utility function values for small input sets
Our remedy is very simple. We set a threshold for input size, denoted by $B$, such that $U(S)$ is always well-defined and meaningful for $|S|\geq B$. For example, in a linear regression with $p$ predictors, observing that $U(S)$ is undefined for $|S|<p$, we may set $B=cp$ for some constant $c\geq 1$. If $|S|<B$, we inject B|S|B-|S| non-informative data points, to elevate the input size to $B$; otherwise no change is made to $U(S)$. As a concrete example for the non-informative distribution, consider a supervised learning scenario with data $D=\{(x_{i},y_{i})\}^{n}_{i=1}$. Here, we can randomly shuffle $y_{i}$’s and use the resulting empirical data distribution as the non-informative distribution $\mathcal{P}_{\rm null}$. Clearly, we can expect that as the size of informative data (i.e., $|S|$) increases, less amount of non-informative data will be injected, and $U(S)$ gradually becomes non-baseline. For further theoretical and algorithmic details, see Appendix C.
@@ 3.5.1 Stochastic Gradient Descent (SGD)
Assumption 3.1 imposes mild smoothness on the utility function $u$, while the remaining assumptions concern the loss function $\ell$. Compared with the classical analysis of SGD stability by Hardt et al. (2016), the only additional requirement is Assumption 3.4, which ensures smoothness of the Hessian. This condition is mild and is satisfied by standard smooth loss functions, such as the squared loss, when the prediction function is twice continuously differentiable in terms of parameters $w$.
@@
$\left|U(\mathcal{S}\cup\{z_{1},z_{1}^{\prime}\})-U(\mathcal{S}\cup\{z_{1},z_{2}\})-U(\mathcal{S}\cup\{z_{1}^{\prime},z_{2}^{\prime}\})+U(\mathcal{S}\cup\{z_{2},z_{2}^{\prime}\})\right|\lesssim s^{-(3/2+\upsilon)}.$
3.5.2 Influence Function (IF)
The influence function (IF) method (Hampel, 1974; Koh and Liang, 2017) considers the following regularized empirical loss:
Like in Section 3.5.1, suppose the utility can be written as $U(S)=u(\widehat{\theta}_{S})$. Under mild conditions (see Proposition 3), standard IF theory (Hampel, 1974; Koh and Liang, 2017) implies that
@@ 3.5.2 Influence Function (IF)
Suppose $u$ is continuously differentiable with LL‑Lipschitz and bounded gradient. Assume that the loss function $\ell(\theta;z)$ is convex in $\theta$, three times continuously differentiable with all derivatives up to order three uniformly bounded. Then for any $z_{1},z_{1}^{\prime},z_{2},z_{2}^{\prime}\in\mathcal{Z}$,
@@ 4 Experiments
We empirically compare our method (FGSV) to the approach of first computing individual Shapley values using SOTA methods and then summing them to form a group valuation. The results demonstrate the superiority of our method in both speed and approximation accuracy and its faithfulness under shell company attack in applications such as copyright attribution in generative AI and explainable AI. We report the key findings in the main paper and relegate full details to Appendix B.
@@ 4.1 Approximation accuracy and computational efficiency
In this experiment, we compare our method to the following benchmarks: (1) Permutation-based estimator (Castro et al., 2009; Ghorbani and Zou, 2019) that averages marginal contributions over random data permutations; (2) Group Testing-based estimator (Jia et al., 2019; Wang and Jia, 2023a) that uses randomized group inclusion tests to estimate pairwise Shapley value differences; (3) Complementary Contribution estimator (Zhang et al., 2023) that uses stratified sampling of complementary coalitions; (4) One-for-All estimator (Li and Yu, 2024b) that uses weighted subsets and reuses utility evaluations; (5) KernelSHAP (Lundberg, 2017), a regression-based method with locally weighted samples; (6) Unbiased KernelSHAP (Covert and Lee, 2021), KernelSHAP with a bias-correction; and (7) LeverageSHAP (Musco and Witter, 2025) that speeds up KernelSHAP by weighted sampling based on leverage scores.
@@
We consider the Sum-of-Unanimity (SOU) cooperative game (Li and Yu, 2024b), where the individual Shapley values have a closed-form expression. Specifically, the utility function is defined as $U(S)=\sum_{j=1}^{d}\alpha_{j}\mathbf{1}_{\mathcal{A}_{j}\subseteq S}$, and the corresponding Shapley value for player $i\in[n]$ is given by $\operatorname{SV}(i)=\sum_{j=1}^{d}\frac{\alpha_{j}}{|\mathcal{A}_{j}|}\mathbf{1}_{i\in\mathcal{A}_{j}}$. Here, each subset $\mathcal{A}_{j}$ is generated by sampling a random size uniformly from $1$ to $n$ and then drawing that many players without replacement. We set $d=n^{2}$ and the coefficient $\alpha_{j}$ as the average of the weights of all players $i\in\mathcal{A}_{j}$, where each player $i$ has a weight $\frac{(i\bmod 4)}{4}$. We partition all players into $4$ groups based on their indices: $S_{k}=\big\{i\in[n]:i\equiv k-1(\bmod 4)\big\}$ for $k\in[4]$, and our goal is to estimate the group value $\operatorname{FGSV}(S_{k})=\sum_{i\in S_{k}}\operatorname{SV}(i)$. Given a fixed budget of 20,000 utility function evaluations, we record the absolute relative error of the $\operatorname{FGSV}$ estimate every 200 iterations for each method. We summarize convergence behavior using the Area Under the Convergence Curve (AUCC), defined as: $\operatorname{AUCC}(S_{k})=\frac{1}{100}\sum_{t=1}^{100}|\frac{\operatorname{FGSV}(S_{k})-\widehat{\operatorname{FGSV}}^{(200\cdot t)}(S_{k})}{\operatorname{FGSV}(S_{k})}|,$ which captures both speed and stability of convergence. We also report the Absolute Relative Error (ARE) of the final estimate: $\operatorname{ARE}(S_{k})=\left|\frac{\operatorname{FGSV}(S_{k})-\widehat{\operatorname{FGSV}}^{(20000)}(S_{k})}{\operatorname{FGSV}(S_{k})}\right|.$ Lower AUCC and ARE indicate faster convergence and better accuracy, respectively. We record the average runtime per iteration for each method.
Figure 2: Performance comparison in the SOU game. Top: Our method (FGSV) achieves the lowest AUCC and ARE across all problem sizes. Bottom: Our method costs the lowest runtime per iteration.